Tangents to the ellipse 16x2+9y2=144 making equal intercepts on both the axes are. It is a multiple choice question.
 

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
You should also gives options So please repost this question or you can also attach an image. I will be happy to help you.
supraja venkatraman
5 years ago
The options for the above posted question is 
A) y=x+5
B) y=x-5
C) y= -x+5 
D) y= -x-5
Please answer the question. I got the answer to be c and d as the slope has to be -1 to have equal intercept but the answer key is given to be a b c d. Is the key wrong?